# PAX-TS: Model-agnostic multi-granular explanations for time series forecasting via localized perturbations

### v1

@@ IV-E Low granularity: Multivariate cross-channel correlations
IV-F Comparison to ShapTime and TS-MULE
(a) ShapTime
(b) TS-MULE
We compare the explanations generated by PAX-TS with two explainability algorithms for forecasting: TS-MULE [10] and ShapTime [11]. Fig. 7 shows explanations as generated by the two methods, following the visualization style of the original papers, for the same sample of the CIF dataset previously shown in Fig. 3, where we highlighted high-granularity time step importance in relation to the maximum of the forecast.
Both ShapTime and TS-MULE are not able to capture importance for specific properties of interest, such as the maximum of the forecast, while PAX-TS can derive importance scores for different properties of interest, including extrema, summary statistics, and values at a given index. Further, TS-MULE’s results are always based on segment importance, rather than timestep importance, having a lower granularity than PAX-TS. The granularity of ShapTime can be configured with the number of segments, up to the time step level. However, as ShapTime evaluates all possible distinct subsets, this results in a time complexity of $O(2^{b})$ inference operations, compared to the $O(|\mathcal{A}|b)$ inference operations of PAX-TS to generate importance scores at the time-step-level.
Further, given a trained model, both ShapTime and PAX-TS are deterministic, while TS-MULE produces random results with high variance, leading to high instability in the outcomes. Since ShapTime observes changes in the mean of the forecast, the resulting importance scores are similar to PAX-TS, when setting $\pi(\cdot)=Mean$.
Both ShapTime and TS-MULE are not suitable for application to multivariate data, whereas PAX-TS can be used to analyze cross-channel correlations, which produces a correlation graph, as shown in Fig. 5. Additionally, the time step correlation heatmaps of PAX-TS are able to capture temporal dependency patterns in the data, which can not be highlighted with either of the competitor explainability algorithms.
V Discussion
PAX-TS can generate explanations for time series forecasting models at different granularities, which are suitable for different use cases, depending on the aim of the end-user. In Sec. I, we introduced typical questions that arise from end-users, which we relate to PAX-TS’ explanations here.
“Why is the forecast of productivity at 15:00 so low?”
“How can we increase the trend of the forecast?”
Lastly, PAX-TS can answer questions on cross-channel correlations of multivariate data with a graph, as shown with the example in Fig. 5. For more specific queries between channels, time-step-level index correlation can be used to illustrate the relation between the channels on a higher granularity.
To apply PAX-TS, a top-down approach with increasingly higher explanation granularity can be employed: For multivariate data, first, a cross-channel correlation graph can be used to unveil correlations between different channels of the data. With this information, a heatmap of index correlations can be visualized to illustrate either the correlation between two channels or of a channel with itself at a higher granularity. Finally, when the end-user is investigating a specific time step, extremum, or summary statistic, a stemplot can be generated to assess the time step importance of the specified sample for the given property of interest. With this highest level of granularity, PAX-TS can assist the end-user in comprehending time-step-level importance of the data.
PAX-TS can be applied in a diverse range of application scenarios: In the agriculture domain, PAX-TS can be used to explain soil moisture forecasts [29], where farmers are aiming to optimize their irrigation system. With our method, the system can be tuned to, for example, increase the minimum moisture level. In the context of a smart building digital twin [30], the method can provide detailed explanations for CO2 concentration forecasts to, for example, a facility manager, highlighting how different rooms in the building contribute to the overall CO2 levels. The work on hospital occupancy forecasts by Avinash et al. [31] could be extended with PAX-TS to make use of state-of-the-art forecasting methods while providing post-hoc explanations for hospital managers.
VI Conclusion
In this paper, we introduced PAX-TS to address the lack of suitable explainability methods for time series forecasting. PAX-TS is model-agnostic and post-hoc applicable to any time series forecasting model, and it operates based on localized perturbations of the input data. We found that the explanations provided by PAX-TS are both human-comprehensible and suitable to address practical questions on time series forecasts from end-users. In comparison with ShapTime and TS-MULE, two state-of-the-art XAI methods, PAX-TS provides both higher and lower granularity explanations. Further, the method is applicable to multivariate data, where it can highlight cross-channel correlations as well as temporal dependency patterns. In our experiments, we classified the identified temporal dependency patterns into six distinct classes. Across all datasets and algorithms, we found that diagonal and bipolar regional dependency patterns tend to perform best. We believe that future work can apply PAX-TS for different algorithms on other datasets, which will contribute to our understanding of temporal dependency patterns.
@@ VI Conclusion
TABLE IV: Errors of the evaluated algorithms on each dataset and corresponding temporal dependency pattern. (A) Diagonals, (B) - Diagonals (End), (C) - Last-Timestep, (D) - Bipolar Regions, (E) - Fully correlated, (F) - Other. The best-performing algorithm for each metric is highlighted in bold. p-Values are calculated using the Bonferroni-Dunn test compared to the Naïve model, with significance ($p<0.05$) highlighted in bold.

### v2

@@ Appendix D Multivariate cross-channel correlations
Fig. 8 visualizes an input sample ($\mathbf{x}$) as well as a perturbed sample ($\mathbf{x}^{\prime}$) of channel 5 of the ETTh1 dataset. The second row shows the corresponding forecasts of MultiPatchFormer for both samples ($\mathbf{\hat{y}}$$\mathbf{\hat{y}}^{\prime}$) in channel 0. This demonstrates an example where the model does not consider one channel to make a prediction for another one.
Figure 8: Original and perturbed sample from channel 5 of ETTh1 (first row), and forecast of channel 0 for both samples (second row), as produced by MultiPatchFormer.
Appendix E PAX-TS for long-term forecasting
We investigated PAX-TS’ capability for explaining forecasting models on long-term forecasts, which is a common evaluation approach in the literature (see e.g. [zhou2021informer, liu2023itransformer, wang2024timemixer]). Fig. 9 demonstrates an example of a time step correlation matrix for iTransformer on the M4 Hourly dataset, produced by PAX-TS. The matrix shows the bipolar regions pattern, across a long-term horizon, reflected by the positive and negative change ratios with clear structure. This illustrates the applicability of our method to time series forecasting setups of any combination of input window length $b$ and forecasting horizon $h$. This further demonstrates that all methods and visualizations presented in this paper are applicable to both short-term and long-term forecasting setups.
Figure 9: Time step correlation matrix, as produced by PAX-TS, for long-term forecasting, with $b=96$, and $h=256$. iTransformer is used as the forecasting model on the dataset M4 Hourly.
Appendix F Forecasting error and temporal dependency patterns
Table VI gives a detailed overview of all evaluated forecasting models’ performance on each dataset. The table further highlights which temporal dependency pattern each model showed when analyzing a time step correlation matrix, such as the ones shown in Fig. 3. The last row of Table VI shows the mean rank of each model across all datasets as well as the p-value, showing whether the performance is significantly different from the naïve forecasting model based on the Bonferroni-Dunn test. Notably, TimeMixer, MultiPatchFormer, and iTransformer perform best, with a significant Bonferroni-Dunn test [dunn1961multiple] compared to the Naïve model. On the multivariate datasets Covid, Exchange Rate, and Illness, none of the benchmarked models achieves considerably better performance than the naïve forecaster, highlighting the difficulty of the multivariate forecasting problem. We additionally measured the total inference time across all datasets (Algorithm 1 line (2) + line (6)), where the LLM TimesFM incurred a significantly higher time (1390s) compared to TimeMixer (20.57s), MultiPatchFormer (8.89s), and iTransformer (1.08s).
@@ Appendix F Forecasting error and temporal dependency patterns
TABLE VI: Errors of the evaluated algorithms on each dataset and corresponding temporal dependency pattern. (A) Diagonals, (B) - Diagonals (End), (C) - Last-Timestep, (D) - Bipolar Regions, (E) - Fully correlated, (F) - Other. The best-performing algorithm for each metric is highlighted in bold. p-Values are calculated using the Bonferroni-Dunn test compared to the Naïve model, with significance ($p<0.05$) highlighted in bold.